OpenStudy (anonymous):

How do you factor t^2-10t+21?

OpenStudy (anonymous):

It's \[ (t-3)(t-7) \]

OpenStudy (anonymous):

(t - 3)·(t - 7)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

factorizes as well: a number that I added or subtracted as a result of the number of medium and found the same numbers to be multiplied as a result give me the expression that in this case was 21 signs depend on the original expression a= -3 b= -7 -7-3= 10 -7*-3=21 in this case= (t-a ) (t-b)